From the top of a tower first body thrown up and reaches the ground in 9 s. Second body thrown down with the same initial speed and reaches ground in 4 s, then time taken by freely falling third body to reach to the ground from same height is


1
Expert's answer
2019-09-02T15:43:33-0400

Displacement of a first body (thrown up)


"d=-v_it_1+\frac{gt_1^2}{2}"

Displacement of a second body (thrown down)


"d=v_it_2+\frac{gt_2^2}{2}"

So


"-v_it_1+\frac{gt_1^2}{2}=v_it_2+\frac{gt_2^2}{2}"

Therefore, the initial velocity


"v_i=\frac{g}{2}(t_1-t_2)=\frac{10}{2}(9-4)=25\:\rm{m/s}"

The height of a tower


"d=-25\times 9+\frac{10\times 9^2}{2}=180\:\rm{m}"

Finally, time taken by freely falling third body


"t_3=\sqrt{\frac{2d}{g}}=\sqrt{\frac{2\times 180}{10}}=6\:\rm{s}"